OpenStudy (anonymous):

solve or simplify \[\sqrt[6]{128^{7}}\div \sqrt[6]{2x}\]

OpenStudy (anonymous):

\[\sqrt[6]{(128^{7})}\div \sqrt[6]{2x}\] \[\sqrt[6]{(128^{6})*128\div(2x)}\] \[\sqrt[6]{(128^{6})*64\div(x)}\] \[\sqrt[6]{(128^{6})*(2^{6})\div(x)}\] \[(128*2)\div(\sqrt[6]{x})= 256\div(\sqrt[6]{x})\] You can also write it as : \[(256 \times \sqrt[6]{x})\div x^{6}\] Voila

OpenStudy (anonymous):

\[9\sqrt12t^3+4t \sqrt128t-2t \sqrt48t\]